A flywheel, together with its shaft , has a total mass of 300kg, and its radius of gyration is 900mm. If the effect of bearing friction is equivalent to a couple of 70Nm, calculate the breaking torque required to bring the flywheel to rest from a speed of 12rev/s in 8s.
1
Expert's answer
2020-11-10T06:52:56-0500

Total mass "M=300kg"

Radius of gyration "k=900mm"

Coupled form by friction of bearing "C=70Nm"


Moment of inertia "I=MK^2"


"I=300/times (900)^2\\=243kgm^2"


Speed "\omega=12rev/s"


Work done by shaft "W=\dfrac{1}{2}I\omega^2"


"\dfrac{1}{2}\times 243\times (12)^2\\=243\times 72=17496Nm"


Hence the breaking torques required to brake the shaft="17496-70=17426Nm"